# Controllable Prompt Tuning For

### v1

@@ 5 Experiments
In this section, we evaluate the effectiveness of the proposed CPT method on benchmark image datasets in the presence of spurious features: Waterbirds [68], CelebA [56], MetaShift [52] and ISIC [15]. Due to the space limit constraint, we briefly provide an overview of our experimental setup below, see Appendix B for more details and additional results.
5.1 Datasets
Waterbirds is created by placing bird photos from the Caltech-UCSD Birds dataset [78] with background images taken from Places [91]. The target attributes are foreground objects (i.e. landbird or waterbird) while spurious correlations are background (i.e. land or water landscape).
[…]
MetaShift is another real-world dataset [52] for cat-dog classification, where the cat and dog images are spuriously correlated with indoor (e.g. bed) and outdoor objects (e.g. bike). For evaluation, models are given dog and cat images associated with shelf background, which does not appear in the training set. Thus, this dataset represents a more coherent distributional shift, compared to other benchmarks.
ISIC is an even more challenging dataset with the appearance of multiple spurious features [15], which includes dermoscopic images of skin lesions with target attributes benign or melanoma. Visual examples of those datasets, along with the size for each group in train, validation, and test sets are given in Appendix B.
5.2 Baselines
Following the previous studies, we compare CPT against other state-of-the-art methods for mitigating spurious correlations, including GroupDRO [68], Deep Feature Reweighting (DFR) [42], Progressive Data Expansion (PDE) [19], Subsample [19], Common Gradient Descent (CGD) [64] with the access to the group information during training. Besides, we consider those methods require group labels for validation: Learning from failure (LfF) [60], Just Train Twice (JTT) [53], Predictive Group Invariance (PGI) [3], Contrastive Input Morphing (CIM) [75], Correct-n-Contrast (CNC) [90], or those do not use group annotations at all like Empirical Risk Minimization (ERM) and Automatic Feature Reweighting (AFR) [65]. Moreover, we also report the results from two recent studies of spurious correlation on ViT [29] and CLIP [85].
Regarding the baselines for those experiments further taking the distribution shift into account, we follow exactly the protocol of [82] and compare our method to Empirical Risk Minimization (ERM), Upweighting (UW), Invariant risk minimization (IRM) [6], Information Bottleneck based Invariant Risk Minimization (IB-IRM) [4], Variance Risk Extrapolation (V-REx) [44], CORrelation ALignment (CORAL) [73], and Fish [71]; instance reweighting methods: GroupDRO [68], JTT [53], Domain mixup (DM-ADA) [83], Selective Augmentation (LISA) [86] and the current state-of-the-art method on those benchmarks Discover and Cure (DISC) [82].
5.3 Model training and evaluation
We examine our proposed method on ResNet50 [35], ViT-B/16 [23] and different variants of OpenAI’s CLIP [66] to align with previous work [42, 29, 85]. For prompt tuning on ViT, we employ prompts of lengths $L=5$ in the MetaShift and ISIC, or $L=10$ in the Waterbirds and CelebA datasets. Similarly, $L$ is set to $16$ for all CLIP backbones throughout our experiment. Unless otherwise explicitly mentioned in the experiment, the value of the controlling vector $c$ is set to the default value of $(1,\dots,1)$. Other detailed configurations are relegated to Appendix B.
Due to the large number of groups in ISIC, we calculate the AUROC score [10], following [82, 9] while computing average, worst-group and mean (if possible) accuracy scores for other datasets. For a fair and precise comparison, results of baselines are taken from recent studies [19, 82, 64, 90, 41] and reported directly from original papers.
5.4 Results
Waterbirds and CelebA We start by validating CPT on two well-established benchmarks for investigating spurious correlation, namely Waterbirds and CelebA in Table 5.4. From the results, we see that CPT consistently achieves the best worst-group accuracy compared with the recent state-of-the-art methods such as DFR and PDE on both datasets. Remarkably, our method only updates a tiny portion ($0.4\%$) of parameters. As such, CPT is much more computationally efficient than methods that need to update the entire network. Furthermore, we see that CPT has strong gains in average accuracy, thus not only outperforming other spurious correlation methods on both metrics but also closing the gap to the average accuracy of ERM. Another interesting observation is that those baselines aim to balance the learning between groups via downsampling the minority groups (Subsample), reweighting group losses inversely proportional to the group populations (UW), or computing scaling weights using group training performance to upweight the worst-group loss (GroupDRO), are effective in alleviating spurious correlation. Even though, CPT still exhibits impressive performance compared to the aforementioned methods by using an adaptive balancing mechanism at each iteration instead of relying on fixed reweighting coefficients.
@@ 5.4 Results
De-bias ViT and CLIP Motivated by recent studies focusing on investigating spurious correlations of pre-trained transformer-based models [85, 18], we apply our proposed algorithm on ViT and CLIP and then compare them against those work. It is worth noting that, the authors did not experiment on CelebA with the blond hair target. Hence, we only show their original results on Waterbirds in Figure 3 and Table 2. As can be seen in Figure 3, while CPT enjoys much less computation and memory overhead (only $0.1\%$ and $0.04\%$ compared to ViT-B/16 itself and BiT-M-R50x3 [29]) it still outperforms the second-best model by a large margin. This significant improvement comes from our prompting design and optimization contribution, respectively. Details of backbones are given in Appendix B.
@@
[…]

### v2

@@ 5 Experiments
In this section, we evaluate the effectiveness of the proposed CPT method on benchmark image datasets in the presence of spurious features: Waterbirds (Sagawa et al., 2019), CelebA (Liu et al., 2015), MetaShift (Liang & Zou, 2021) and ISIC (Codella et al., 2019). Due to the space limit constraint, we briefly provide an overview of our experimental setup below, see Appendix B for more details and additional results. Our implementation is available at https://github.com/VietHoang1512/CPT.
@@
5.1 Datasets
[…]
MetaShift is another real-world dataset for cat-dog classification, where the cat and dog images are spuriously correlated with indoor (e.g. bed) and outdoor objects (e.g. bike). For evaluation, models are given dog and cat images associated with shelf backgrounds, which do not appear in the training set. Thus, this dataset represents a more coherent distributional shift, compared to other benchmarks. ISIC is an even more challenging dataset with the appearance of multiple spurious features (Codella et al., 2019), which includes dermoscopic images of skin lesions with target attributes benign or melanoma. Visual examples of those datasets, along with the size for each group in train, validation, and test sets are given in Appendix B.
5.2 Baselines
Following the previous studies, we compare CPT against other state-of-the-art methods for mitigating spurious correlations, including DFR (Kirichenko et al., 2022), GroupDRO (Sagawa et al., 2019), PDE (Deng et al., 2023), Subsample (Deng et al., 2023), CGD (Piratla et al., 2021) with the access to the group information during training. Besides, we consider those methods require group labels for validation: LfF (Nam et al., 2020), JTT (Liu et al., 2021b), PGI (Ahmed et al., 2020), CIM (Taghanaki et al., 2021), CNC (Zhang et al., 2022), or those do not use group annotations at all like ERM and AFR (Qiu et al., 2023). Moreover, we also report the results from two recent studies of spurious correlation on ViT (Ghosal & Li, 2023) and CLIP (Yang et al., 2023b; Dehdashtian & Boddeti, 2024).
Regarding the baselines for those experiments further taking the distribution shift into account, we follow exactly the protocol of Wu et al. (2023b) and compare our method to Upweighting (UW), IRM (Arjovsky et al., 2019), IB-IRM (Ahuja et al., 2021), V-REx (Krueger et al., 2021), CORAL (Sun et al., 2017), and Fish (Shi et al., 2021); instance reweighting methods: JTT (Liu et al., 2021b), DM-ADA (Xu et al., 2020), LISA (Yao et al., 2022) and the current state-of-the-art method, DISC (Wu et al., 2023b).
5.3 Model training and evaluation
We examine our proposed method on ResNet50 (He et al., 2016), ViT-B/16 (Dosovitskiy et al., 2020) and different variants of OpenAI’s CLIP (Radford et al., 2021) to align with previous work (Kirichenko et al., 2022; Ghosal & Li, 2023; Yang et al., 2023b). For prompt tuning on ViT, we employ prompts of lengths $L=5$ in the MetaShift and ISIC, or $L=10$ in the Waterbirds and CelebA datasets. Similarly, $L$ is set to $16$ for all CLIP backbones throughout our experiment. Unless otherwise explicitly mentioned in the experiment, the value of the controlling vector $c$ is set to the default value of $(1,\dots,1)$. Other detailed configurations are relegated to Appendix B.
Due to the large number of groups in ISIC, we calculate the AUROC score (Bradley, 1997), following (Wu et al., 2023b; Bissoto et al., 2020) while computing average, worst-group and mean (if possible) accuracy scores for other datasets. For a fair and precise comparison, results of baselines are taken from recent studies (Deng et al., 2023; Wu et al., 2023b; Piratla et al., 2021; Zhang et al., 2022; Kim et al., 2023) and reported directly from original papers.
5.4 Results
Waterbirds and CelebA. We start by validating CPT on two well-established benchmarks for investigating spurious correlation, namely Waterbirds and CelebA in Table 5. From the results, we see that CPT consistently achieves the best worst-group accuracy compared with the recent state-of-the-art methods such as DFR and PDE on both datasets. Remarkably, our method only updates a tiny portion ($0.4\%$) of parameters. As such, CPT is much more computationally efficient than methods that need to update the entire network. Furthermore, we see that CPT has strong gains in average accuracy, thus not only outperforming other spurious correlation methods on both metrics but also closing the gap to the average accuracy of ERM. Another interesting observation is that those baselines aim to balance the learning between groups are effective in alleviating spurious correlation, e.g. via downsampling the majority groups (Subsample), reweighting group losses inversely proportional to the group populations (UW), or computing scaling weights using group training performance to upweight the worst-group loss (GroupDRO). Even though, CPT still exhibits impressive performance compared to the aforementioned methods by using an adaptive balancing mechanism at each iteration instead of relying on fixed reweighting coefficients.
@@ 5.4 Results
De-bias ViT and CLIP. Motivated by recent studies focusing on investigating spurious correlations of pre-trained transformer-based models (Yang et al., 2023b; Dehdashtian & Boddeti, 2024), we apply our proposed algorithm on ViT and CLIP and then compare them against those work. It is worth noting that, the authors did not experiment on CelebA with the blond hair target. Hence, we only show their original results on Waterbirds in Figure 3 and Table 2. As can be seen in Figure 3, while CPT enjoys much less computation and memory overhead (only $0.1\%$ and $0.04\%$ compared to ViT-B/16 itself and BiT-M-R50x3 (Ghosal & Li, 2023)) it still outperforms the second-best model by a large margin. This significant improvement comes from our prompting design and optimization contribution, respectively. Details of backbones are given in Appendix B.